OpenStudy (anonymous):

How many units apart are the points (4, -3) and (-4, 3)?

OpenStudy (anonymous):

Use the distance formula

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ -3}})\quad &({\color{red}{ -4}}\quad ,&{\color{blue}{ 3}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

64+81 seems a lot for distance.

OpenStudy (anonymous):

\[ \sqrt {(4-(-4))^2 + (3-(-3))^2 }=\sqrt{ 64+36} \]

OpenStudy (anonymous):

It is \[ \sqrt{100}=10 \]

OpenStudy (jdoe0001):

notice \(\bf \implies \sqrt{\qquad }\Longleftarrow\)

OpenStudy (anonymous):

Oh ok. So it would be 10 units apart?